Reports 1-1 of 1 Clear search Modify search
AdV-PSL (High Power fiber amplifier)
pillant, cleva, derossi, chiummo, gosselin, TCS - 11:45 Wednesday 13 March 2019 (45246) Print this report
Laser Amplifier chiller failure

Since a while we noticed a temperature increase of the laser diodes of the power amplifier.
During the last days (see plot 1) the temperature rised close to the high temperature limit (35 °C).
So we checked yesterday the chiller and after the first test we noticed a problem with the water cooling flow.
After some attempts to find the blocked part, the chiller stopped working properly and we found half of the water tank completly freezed (see picture 1),
that was not the case at the begining of the investigation (picture 2).
We swapped the chiller with a spare kindly loaned by the TCS team.
Even if the temperature is now stable with a safe margin, we plan to plug the laser amplifier to the main laser chiller.

 

 

Images attached to this report
Comments to this report:
cleva - 12:00 Wednesday 13 March 2019 (45247) Print this report

The ice was not visible upon our arrival in the EERoom but it may be that the bottom, non visible, part of the reservoir ( 50%) was ~ frozen before our arrival. The ice cube could have settled gradually over the last year and explain the continuous drop of the chiller perf (see plot 2)

Rem.: The reservoir is only ~ 2 l (see picture 1, grey box on the upper / left side) and such a small volume is a weak point of that chiller.

Images attached to this comment
Search
×

Warning

×